ECO 2013                                                                                          Instructor: Zhijing Teng

Assignment 6                                                                                     Due Date:

This assignment is worth 8 points (Each question is worth 0.5 point). No late homework will be accepted without a documented excuse.

Chile and Argentina

Suppose that a typical labor in Chile and Argentina has daily productivities of making steak and wine as following.

Daily Production Possibilities

Steak

Wine

Chile

4

12

Argentina

2

5

Question 1:Chile’s opportunity cost of producing 1 steak is __3__ wine.

Answer:

give up / gain =  

Question 2:Chile’s opportunity cost of producing 1 wine is __0__ steak.

Answer:

Give up / gain =  

Question 3:Argentina’s opportunity cost of producing 1 steak is _2__ wine.

Answer:

give up / gain =  

Question 4:Argentina’s opportunity cost of producing 1 wine is __0__ steak.

Answer:

Give up / gain =  

Smoothie Bar

In an hour, Liz can produce either 30 smoothies or 30 salads. In an hour, Joe can produce either 6 smoothies or 30 salads.

Question 1:Liz’s opportunity cost of producing 1 smoothie is ___ salad.

Answer:

give up / gain =  

Question 2:Liz’s opportunity cost of producing 1 salad is ____ smoothie.

Answer:

Give up / gain =  

Question 3:Joe’s opportunity cost of producing 1 smoothie is ___ salad.

Answer:

give up / gain =  

Question 4:Joe’s opportunity cost of producing 1 salad is ____ smoothie.

Answer:

Give up / gain =  

Bake Sale

On a normal day, David can make 12 cakes or 6 pies. Loren can make 8 cakes or 6 pies.

Question 1:David’s opportunity cost of producing 1 cake is ___ pies.

Answer:

give up / gain =  

Question 2:David’s opportunity cost of producing 1 pie is ____ cakes.

Answer:

Give up / gain =  

Question 3:Loren’s opportunity cost of producing 1 cake is ___ pies.

Answer:

give up / gain =  

Question 4:Loren’s opportunity cost of producing 1 pie is ____ cakes.

Answer:

Give up / gain =  

Jake and Georgie

In a week, Jake can catch 15 fish or pick 30 pineapples. Georgie can catch 20 fish or pick 50 pineapples.

Weekly Production Possibilities

Fish

Pineapples

Jake

15

30

Georgie

20

50

Question 1:Jake’s opportunity cost of getting 1 fish is ___ pineapples.

Answer:

give up / gain =  

Question 2:Jake’s opportunity cost of getting 1 pineapple is ____ fish.

Answer:

Give up / gain =  

Question 3:Georgie’s opportunity cost of getting 1 fish is ___ pineapples.

Answer:

give up / gain =  

Question 4:Georgie’s opportunity cost of getting 1 pineapple is ____ fish.

Answer:

Give up / gain =  
Answer #1

As per rules and time constraint, only one question is allowed. However, I am answering first two. Kindly ask other questions separately. Thank you.

Ans) Opportunity cost is the cost of something that is given up.

1) For Chile÷

Opportunity cost of steak = wine sacrificed ÷ steak gained. = 12÷4=3 wines

Opportunity cost of wine = steak sacrificed ÷ wine gained = 4÷12 = 1/3 or 0.33 steak.

For Argentina ÷

Opportunity cost of steak = wine sacrificed ÷ steak gained = 5÷2 = 5/2 or 2.5 wine

Opportunity cost of wine= steak sacrificed ÷ wine gained = 2÷5= 2/5 or 0.4 steak

2)

Smoothie Salad
Liz 30 30
Joe 6 30

For Liz÷

Opportunity cost of smoothie= salad sacrificed ÷ smoothie gained = 30÷30 = 1 salad

Opportunity cost of salad = smoothie sacrificed ÷ salad gained = 30÷30 = 1 smoothie.

For Joe÷

Opportunity cost of smoothie = salad sacrificed ÷ smoothie gained = 30÷6 = 5salad.

Opportunity cost of salad= smoothie sacrificed ÷ salad gained = 6÷30 = 6/30 or 0.2 smoothie.